How do you solve for linear equations with this problem: 4x+3y=5
-3x-6y=0

Answer by glitzgirl_14(11)

The easiest way to do this problem is to use elimination. So take your first equation and multiply it my 2:
8x+6y=5
-3x-6y=0
Add the x's together:
5x=5
x=1
Put 1 in for x in the second equation:
-3*1-6y=0
-3-6y=0
-6y=3
y=-1/2